# Visual and Textual Prompts for Enhancing Emotion Recognition in Video

### v1

@@ I Introduction
Emotion recognition is a key area in computer vision [1], enabling systems to understand and respond to human emotions in applications like virtual reality, HCI [2], surveillance [3], and mental health [4]. Understanding emotions involves not only analyzing facial expressions but also interpreting complex signals from body language and contextual interactions. Traditional methods for emotion recognition [5, 6, 7, 8] primarily focus on facial features, using handcrafted features or shallow machine learning models [9]. While these approaches have shown success in controlled environments, they often fail in real-world settings where emotions are influenced by intricate interpersonal dynamics, environmental factors, and cultural variations. Recent advances in deep learning and VLLMs have opened new possibilities for emotion recognition. Vision Large Language Models (VLLMs) have emerged in this domain by combining the generalization capabilities of language models with visual understanding. Inspired by the breakthroughs of LLMs such as LLaMA [10], PaLM [11], and GPT-4 [12], VLLMs extend these capabilities to the visual domain through training on large-scale video-text datasets. These models excel in tasks such as visual reasoning [13], scene understanding [14], and object recognition [15], showcasing their ability to learn from complex multi-modal inputs. However, applying VLLMs to emotion recognition remains challenging due to their reliance on nuanced visual cues, including inter-personal dynamics, spatial context, and subtle facial features and the need to train transformer-based models [16] or CLIP-based VLLMs [17] to enhance in-context emotion classification.
@@
Recent studies [18, 19, 20] have explored zero-shot prompting techniques to enhance image-language tasks by guiding Large Language Models (LLMs) with specific visual markers. For example, RedCircle [21] employs red circles around objects to focus a Vision-Language Model’s attention on specific regions, improving tasks like zero-shot referring expression comprehension and keypoint localization. Similarly, ReCLIP [22] uses a region-scoring approach that combines colorful boxes for cropping and blurring irrelevant areas to enhance performance, significantly reducing the gap between zero-shot and supervised models on both real-world and synthetic datasets. Yang et al. [18] introduced fine-grained visual prompting with segmentation masks and a ‘Blur Reverse Mask’ technique that blurs regions outside the target area to preserve spatial context and reduce distractions. While these advances have demonstrated success in tasks like object grounding and segmentation, they face limitations when applied to emotion recognition. Methods like RedCircle and ReCLIP rely heavily on cropping techniques, which can result in the loss of critical contextual information essential for understanding emotions. Similarly, fine-grained visual prompts using Blur Reverse Mask can obscure facial features, introduce ambiguities, and hinder the accurate interpretation of nuanced emotional expressions. Although these methods hold considerable promise, their effectiveness in the domain of emotion recognition has yet to be systematically explored.
@@
  • The paper proposes a novel Set-of-Vision-Text Prompting (SoVTP) approach that integrates spatial information, such as numbered labels, bounding boxes, facial landmarks, and action units, into zero-shot emotion recognition tasks. This method enhances model interpretability and precision by preserving background context while focusing on the individual’s facial region.
  • The SoVTP framework combines multiple modalities—muscle movement analysis (via Action Units), body language assessment, context extraction, and interpersonal emotion tracking—to provide a comprehensive understanding of emotions. This layered visual prompting strategy improves the accuracy of facial emotion recognition, especially in complex social scenes.
  • We present comprehensive experimental results that demonstrate SoVTP significantly improves VLLMs’ emotion recognition accuracy, achieving state-of-the-art performance with substantial gains over existing visual prompting approaches.
@@ II-A Vision Large Language models
Recent advances in VLLMs have been inspired by the success of traditional LLMs such as LLaMA [10], PaLM [11] and GPT-4 [12], which have shown remarkable abilities in zero-shot generalization in a range of NLP tasks. These LLMs have demonstrated the ability to understand complex instructions and perform diverse tasks without explicit fine-tuning. The success of these models has motivated the development of VLLMs, which extend these capabilities to the visual domain by using large-scale video-text paired datasets. VideoLLaMA2 [23] employs an image-level CLIP encoder as the vision backbone with a novel Spatial-Temporal Convolution Connector (STC Connector) for enhanced spatial-temporal representation learning, allowing effective aggregation of frame features while preserving local details without generating excessive video tokens. Video-LLaVA [24] unifies visual representations into the language feature space to create a robust VLLM, achieving significant improvements over existing models on both video and image benchmarks by leveraging mixed datasets for mutual enhancement of image and video understanding. Despite these advancements, existing VLLMs still require substantial fine-tuning of their visual and text encoders to adapt to new tasks, especially those involving nuanced visual content like emotion recognition. In contrast, our work proposes a novel zero-shot architecture for emotion recognition, aimed at using the inherent generalization abilities of VLLMs without additional training.
@@ II-B Prompting methods
Prompt engineering is a commonly used approach in NLP [25, 26]. ToT [27] is a novel inference framework that extends the Chain of Thought approach by enabling language models to explore multiple reasoning paths, make deliberate decisions, and evaluate alternatives, significantly enhancing performance in tasks requiring strategic planning or exploration, such as Creative Writing. Although prompts for large language models have been extensively explored, prompts for vision tasks have received less attention and investigation. Liu et al. [28] addresses limitations in current prompt tuning methods for vision-language models (VLMs), which use a single prompt and struggle to capture diverse visual information. While using multiple prompts could improve diversity, it risks overfitting and imbalances between base and new task performance. AnyRef [29] is a general Multimodal Large Language Model (MLLM) capable of generating pixel-wise object perceptions and natural language descriptions from diverse multi-modality references, employing a refocusing mechanism to improve grounding and referring expression tasks without requiring modality-specific adaptations. FGVP [18] introduces a new zero-shot framework using pixel-level annotations from a generalist segmentation model, leveraging Blur Reverse Mask as an effective visual prompting strategy to enhance instance-level recognition tasks. While these methods have shown promising results in tasks such as semantic segmentation and object grounding, they often struggle with emotion recognition. This challenge arises from their focus on analyzing objects in isolation, neglecting essential global context and detailed facial features necessary for accurately interpreting emotional expressions. To address these shortcomings, our proposed approach emphasizes the extraction of fine-grained facial features across the entire image, maintaining spatial information through the use of spatial annotation such as boxes, numerical labels, facial landmarks, and action units.
@@ III-A Problem Definition
17:    if Check_Overlap($\tilde{F}[k],\tilde{F}[k-1]$):
18:     if Calculate_Area($\tilde{F}[k]$) $>$ Calculate_Area($\tilde{F}[k-1]$):
19:      // Replace the smaller bounding box
20:      visible_faces =$=\neg\tilde{F}[k-1]\land\tilde{F}[k]$
21:   return visible_faces
@@ III-B1 Box detection
$b_{i}=\mathcal{F}(V,\alpha_{i})$ (3)
, where $V$ represents the video sequence, $\alpha_{i}$ denotes the hyperparameters for the RetinaFace model $\mathcal{F}$, and $b_{i}$ corresponds to the bounding box of the $i$-th face.
@@ III-B2 Face Overlap Handling Algorithm
$\centering B_{sorted}=\{b_{1},b_{2},\ldots,b_{n}\}\@add@centering$ (4)
@@
$\begin{split}&\text{Overlap}(\tilde{F}[k],\tilde{F}[j])=\\ &\frac{\text{Area}(\tilde{F}[k]\cap\tilde{F}[j])}{\min(\text{Area}(\tilde{F}[k]),\text{Area}(\tilde{F}[j]))}>\epsilon\end{split}$ (5)
, where $\epsilon$ defines the threshold for significant overlap. If the overlap exceeds this threshold, the algorithm compares the areas of the overlapping faces. The smaller face is discarded, as determined by Eq. (6):
$\tilde{F}[k]=\begin{cases}\hat{F}[k]&\text{if }\text{Area}(\hat{F}[k])>\text{Area}(\hat{F}[j]),\\ \hat{F}[j]&\text{otherwise}.\end{cases}$ (6)
Non-occluded faces are then added to the final set $F_{final}$ using Eq. (7):
$F_{final}\leftarrow F_{final}\cup\{\tilde{F}[k]\}.$ (7)
This process ensures that only faces closer to the camera and free from obstructions are retained. After resolving overlaps and refining the bounding box set $B=\{b_{1},b_{2},\ldots,b_{n}\}$, unique IDs $N=\{1,2,\ldots,n\}$ are assigned to each detected face. These IDs are then used to track faces across the video sequence, enabling distinguishable and interpretable vision prompts for VLLMs. This methodology enhances the model’s ability to interpret complex visual scenes.
@@ III-B3 Muscle Movement Detection and Analysis
$M=L(B)$ (8)
These landmarks are used by the action unit detection model $D$, which outputs the activation scores $A=\{a_{1},a_{2},\dots,a_{m}\}$ for $m$ action units by Eq. (9):
$A=D(\{l_{1},l_{2},\dots,l_{n}\})$ (9)
Once the activation scores are computed, they are ranked to identify the key prominent action units $P_{ac}$, where $P_{ac}$ is a subset of $A=\{a_{1},a_{2},\dots,a_{m}\}$ in Eq. (10):
$P_{ac}=\text{Rank}(A)$ (10)
The ranked action units $P_{ac}$ are then overlaid onto the original face $B$ to generate a visualization $V$ by Eq. (11):
$V=\text{Overlay}(P_{ac},B)$ (11)
@@ III-D Final prompting strategy
The final prompting strategy combines visual and textual elements. Visual Prompts: Overlay target face annotations (numbered bounding box, facial landmarks, AUs) onto the frame, preserving background context (Fig. 2 and Fig. 3). Textual Prompts: Concatenate $T_{\text{context}}$, $T_{\text{body}}$, and $E_{\text{others}}$ into a structured query, e.g.,
$T^{\text{SoVTP}}_{new}=\left[T_{\text{context}};\,T_{\text{body}};\,E_{\text{others}}\right]$ (12)
The enhanced input $(V^{\text{new}},T^{\text{SoVTP}}_{new})$ is fed into the VLLM $\Theta$. The model analyzes and critiques body language, context, and others’ emotions via chain-of-thought reasoning, then outputs the target emotion $A^{o}$ as:
$A^{o}=\Theta(V^{\text{new}},T^{\text{SoVTP}}_{new})$ (13)
@@
IV Experiments
IV-A Models and Settings
IV-B Dataset details

### v3

# Visual and Textual Prompts in VLLMs for Enhancing Emotion Recognition

@@ I Introduction
Emotion recognition is a key area in computer vision [1], enabling systems to understand and respond to human emotions in applications like virtual reality, HCI [2], surveillance [3, 4], and mental health [5]. Robust emotion understanding requires not only analyzing facial expressions but also interpreting body language and contextual interactions—often under challenging visual conditions [6, 7]—and thus benefits from advanced restoration techniques [8]. Traditional methods for emotion recognition [9, 10, 11, 12] primarily focus on facial features, using handcrafted features or shallow machine learning models [13]. While these approaches have shown success in controlled environments, they often fail in real-world settings where emotions are influenced by intricate interpersonal dynamics, environmental factors, and cultural variations.
@@
Recent studies [14, 15, 16] have explored zero-shot prompting techniques to enhance image-language tasks by guiding Large Language Models (LLMs) with specific visual markers. For example, RedCircle [17] employs red circles around objects to focus a Vision-Language Model’s attention on specific regions, improving tasks like zero-shot referring expression comprehension and keypoint localization. Similarly, ReCLIP [18] uses a region-scoring approach that combines colorful boxes for cropping and blurring irrelevant areas to enhance performance. Yang et al. [14] introduced fine-grained visual prompting with segmentation masks and a ‘Blur Reverse Mask’ technique that blurs regions outside the target area to preserve spatial context and reduce distractions. While these advances have demonstrated success in tasks like object grounding and segmentation, they face limitations when applied to emotion recognition. Methods like RedCircle and ReCLIP rely heavily on cropping techniques, which can result in the loss of critical contextual information essential for understanding emotions. Similarly, fine-grained visual prompts using Blur Reverse Mask can obscure facial features, introduce ambiguities, and hinder the accurate interpretation of nuanced emotional expressions. Although these methods hold considerable promise, their effectiveness in the domain of emotion recognition has yet to be systematically explored.
@@
  • The paper proposes a novel Set-of-Vision-Text Prompting (SoVTP) approach that preserves holistic scene context rather than relying on isolated facial-cropping methods by overlaying spatial annotations (e.g., face IDs, landmarks, action units) on full-scene inputs, and integrates body posture, environmental cues, and social dynamics to enrich multimodal signals.
  • The SoVTP framework introduces a multi-stage prompting that treats zero-shot emotion recognition as structured inference over chained reasoning steps, enabling VLLMs to marginalize intermediate evidence without fine-tuning.
  • We present comprehensive experimental results that demonstrate SoVTP significantly improves VLLMs’ emotion recognition accuracy, achieving substantial gains over existing approaches.
@@ II-A Vision Large Language models
Recent advances in VLLMs have been inspired by the success of traditional LLMs such as LLaMA [19], PaLM [20] and GPT-4 [21], which have shown remarkable abilities in zero-shot generalization in a range of NLP tasks. These LLMs have demonstrated the ability to understand complex instructions and perform diverse tasks without explicit fine-tuning. The success of these models has motivated the development of VLLMs, which extend these capabilities to the visual domain by using large-scale video-text paired datasets. VideoLLaMA2 [22] employs an image-level CLIP encoder as the vision backbone with a novel Spatial-Temporal Convolution Connector (STC Connector) for enhanced spatial-temporal representation learning, allowing effective aggregation of frame features while preserving local details without generating excessive video tokens. Video-LLaVA [23] unifies visual representations into the language feature space to create a robust VLLM, achieving significant improvements over existing models on both video and image benchmarks by leveraging mixed datasets for mutual enhancement of image and video understanding. Recent advances in action recognition and affective video analysis have laid important groundwork for emotion recognition: compositional action recognition methods that leverage progressive instance-aware feature learning [24] and hierarchical graph-based cross-inference for group activity recognition [25] demonstrate the value of modeling interactions and temporal dynamics, while skeleton-based techniques such as motion-aware mask feature reconstruction [26, 27] and multi-granularity anchor-contrastive representation learning in semi-supervised settings [28] provide robust strategies for capturing fine-grained body motion cues. In parallel, decade-scale surveys of affective video content analysis [29] and frameworks like CogAware [30], transformer-based models [31] or CLIP-based VLLMs [32] for emotion recognition, social media sentiment integration in domain-specific contexts [33], and unsupervised word-level affect propagation over lexical graphs [34] underscore the importance of combining multimodal signals and knowledge-driven semantics. Furthermore, visual emotion analysis networks—such as object-aroused emotion analysis for images [35] and user-centered approaches integrating emotional states with visual features [36]—highlight effective techniques for emotion recognition. Despite these advancements, existing methods still require substantial fine-tuning of their visual and text encoders to adapt to new tasks, especially those involving nuanced visual content like emotion recognition. In contrast, our work proposes a novel zero-shot architecture for emotion recognition, aimed at using the inherent generalization abilities of VLLMs without additional training.
@@ II-B Prompting methods
Prompt engineering is a commonly used approach in NLP [37, 38, 39]. ToT [40] is a novel inference framework that extends the Chain of Thought approach by enabling language models to explore multiple reasoning paths, make deliberate decisions, and evaluate alternatives, significantly enhancing performance in tasks requiring strategic planning or exploration, such as Creative Writing. Although prompts for large language models have been extensively explored, prompts for vision tasks have received less attention and investigation. Liu et al. [41] addresses limitations in current prompt tuning methods for vision-language models (VLMs), which use a single prompt and struggle to capture diverse visual information. While using multiple prompts could improve diversity, it risks overfitting and imbalances between base and new task performance. AnyRef [42] is a general Multimodal Large Language Model (MLLM) capable of generating pixel-wise object perceptions and natural language descriptions from diverse multi-modality references, employing a refocusing mechanism to improve grounding and referring expression tasks without requiring modality-specific adaptations. FGVP [14] introduces a new zero-shot framework using pixel-level annotations from a generalist segmentation model, leveraging Blur Reverse Mask as an effective visual prompting strategy to enhance instance-level recognition tasks. While these methods have shown promising results in tasks such as semantic segmentation and object grounding, they often struggle with emotion recognition. This challenge arises from their focus on analyzing objects in isolation, neglecting essential global context and detailed facial features necessary for accurately interpreting emotional expressions. To address these shortcomings, our proposed approach emphasizes the extraction of fine-grained facial features across the entire image, maintaining spatial information through the use of spatial annotation such as boxes, numerical labels, facial landmarks, and action units.
@@ III-A Problem Definition
17:    if Check_Overlap($\tilde{F}[k],\tilde{F}[k+1]$):
18:     // Discard the smaller bounding box
19:     visible_faces =$=\neg\tilde{F}[k+1]\land\tilde{F}[k]$
20:   return visible_faces
@@ III-B1 Box detection
$b_{i}=\mathcal{F}(V,\alpha_{i}),$ (3)
where $V$ represents the video sequence, $\alpha_{i}$ denotes the hyperparameters for the RetinaFace model $\mathcal{F}$, and $b_{i}$ corresponds to the bounding box of the $i$-th face.
@@ III-B2 Face Overlap Handling Algorithm
$\centering B_{sorted}=\{b_{1},b_{2},\ldots,b_{n}\},\@add@centering$ (4)
@@
$\begin{split}&\text{Overlap}(\tilde{F}[k],\tilde{F}[j])=\\ &\frac{\text{Area}(\tilde{F}[k]\cap\tilde{F}[j])}{\min(\text{Area}(\tilde{F}[k]),\text{Area}(\tilde{F}[j]))}>\epsilon,\end{split}$ (5)
where $\epsilon$ defines the threshold for significant overlap. $\tilde{F}[k]$ is the current face being considered for inclusion. $\tilde{F}[j]$ is an already accepted face in the final set $F_{final}$. The face $F_{a}$ is set to the current face $\tilde{F}[k]$ only if the overlap with the already accepted face $\tilde{F}[j]$ is less than or equal to $\epsilon$. Otherwise, $F_{a}$ is set to None, meaning the current face is discarded because it is considered occluded or overlaps too much, as determined by Eq. (6):
$F_{a}=\begin{cases}\tilde{F}[k]&\text{if }\text{Overlap}(\tilde{F}[k],\tilde{F}[j])<=\epsilon,\\ None&\text{otherwise}.\end{cases}$ (6)
Then, $F_{a}$ (if not None) is added to the final set $F_{final}$ using Eq. (7):
$F_{final}\leftarrow F_{final}\cup\{F_{a}\}.$ (7)
This process ensures that only larger faces free from obstructions are retained. After resolving overlaps and refining the bounding box set $B=\{b_{1},b_{2},\ldots,b_{n}\}$, unique IDs $N=\{1,2,\ldots,n\}$ are assigned to each detected face. These IDs are then used to track faces across the video sequence, enabling distinguishable and interpretable vision prompts for VLLMs. This methodology enhances the model’s ability to interpret complex visual scenes.
@@ III-B3 Muscle Movement Detection and Analysis
$M=L(B).$ (8)
These landmarks correspond to action units, such as AU1 (Inner Brow Raiser), which are located at facial landmarks (20, 21, 22, 23, 24, 25). The detection model $D$ [44] outputs the activation scores $A=\{a_{1},a_{2},\dots,a_{m}\}$ for $m$ action units as defined in Eq. (9):
$A=D(\{l_{1},l_{2},\dots,l_{n}\}).$ (9)
Once the activation scores are computed, they are ranked to identify the key prominent action units $P_{ac}$, based on the initial dominant emotion extracted by Py-Feat. Here, $P_{ac}$ is a subset of $A=\{a_{1},a_{2},\dots,a_{m}\}$ as defined in Eq. (10):
$P_{ac}=\text{Rank}(A).$ (10)
The ranked action units $P_{ac}$ are then overlaid onto the original face $F$ at the positions of related facial landmarks to generate a visualization $V$ by Eq. (11):
$V=\text{Overlay}(P_{ac},F).$ (11)
@@ III-C2 Combined Vision-Text Prompts
Figure 6: Overview of the multi-stage prompting strategy for emotion recognition using Vision Large Language Models (VLLMs). The approach decomposes the reasoning process into five sequential stages: (1) contextual scene understanding, (2) body language interpretation, (3) inference of emotions from surrounding individuals, (4) detection of facial action units (AUs), and (5) final emotion prediction through self-correction by integrating information from previous stages. Each stage consists of a specific question $Q_{i}$, intermediate contextual reasoning $C_{i}$, and answer $A_{i}$, facilitating a structured chain-of-thought that enhances robustness and accuracy in emotion recognition within complex social environments.
@@ III-D Final prompting strategy
In this section, we introduce preliminaries of reasoning with VLLM prompting. For standard prompting, given the reasoning question $\mathcal{Q}$, prompt $\mathcal{T}$, and a parameterized probabilistic vision large language model $p_{vllm}$, we aim to maximize the likelihood of an answer $\mathcal{A}$:
$p(\mathcal{A}\mid\mathcal{T},\mathcal{Q})=\prod_{i=1}^{|\mathcal{A}|}p_{vllm}\!\bigl{(}a_{i}\,\big{|}\,\mathcal{T},\mathcal{Q},a_{<i}\bigr{)}.$ (12)
where $a_{i}$ is the $i$-th answer token and $|\mathcal{A}|$ is the answer length.
For the proposed multi-stage prompting strategy that involves $K$ reasoning stages, each stage $i$ corresponds to a question-answer pair $(\mathcal{Q}_{i},\mathcal{A}_{i})$. In addition, we incorporate intermediate reasoning steps $\mathcal{C}_{i}$ chain-of-thought at each stage. The overall prompt $\mathcal{T}$ is thus composed as:
$\mathcal{T}=\{(\mathcal{Q}_{i},\mathcal{C}_{i},\mathcal{A}_{i})\}_{i=1}^{K},$
where
  • $\mathcal{Q}_{i}$: the question at stage $i$ (e.g., context, body language, other’s emotions, facial action units, final integration),
  • $\mathcal{C}_{i}$: intermediate reasoning or evidence at stage $i$,
  • $\mathcal{A}_{i}$: the answer at stage $i$.
The final answer likelihood is expressed by marginalizing over all possible reasoning trajectories $\mathcal{C}=\{\mathcal{C}_{i}\}_{i=1}^{K}$:
$p(\mathcal{A}\mid\mathcal{Q})=\sum_{\mathcal{C}}p(\mathcal{A}\mid\mathcal{Q},\mathcal{C})p(\mathcal{C}\mid\mathcal{Q}),$ (13)
with the two factors defined by
$p(\mathcal{C}\mid\mathcal{Q})$ $=\prod_{i=1}^{|\mathcal{C}|}p_{\text{vllm}}(c_{i}\mid\mathcal{Q},c_{<i}),$ (14)
$p(\mathcal{A}\mid\mathcal{Q},\mathcal{C})$ $=\prod_{j=1}^{|\mathcal{A}|}p_{\text{vllm}}(a_{j}\mid\mathcal{Q},\mathcal{C},a_{<j}),$ (15)
where $c_{i}$ denotes the $i$-th step out of the $|\mathcal{C}|$ total reasoning steps.
In the proposed multi-stage SoVTP Prompting, Stage 1 (Context): $\mathcal{Q}_{1}$ asks for scene context; answer $\mathcal{A}_{1}$ summarizes setting.Stage 2 (Body Language): $\mathcal{Q}_{2}$ queries body language; $\mathcal{A}_{2}$ gives nonverbal cues. Stage 3 (Other’s Emotions): $\mathcal{Q}_{3}$ assesses surrounding individuals’ emotions; $\mathcal{A}_{3}$ characterizes social dynamics. Stage 4 (Facial Action Units): $\mathcal{Q}_{4}$ identifies key facial action units; $\mathcal{A}_{4}$ maps to emotional muscle activations. Stage 5 (Self-correction stage): $\mathcal{Q}_{5}$ integrates prior stages via $\mathcal{C}_{i}$ and outputs final emotion $\mathcal{A}_{5}$. The whole process can be described in Fig. 6.
The final prompting strategy combines visual and textual elements. Visual Prompts: Overlay target face annotations (numbered bounding box, facial landmarks, AUs) onto the frame, preserving background context (Fig. 2 and Fig. 3). Textual Prompts include $T_{\text{context}}$, $T_{\text{body}}$, $T_{\text{AU}}$ and $E_{\text{others}}$
$T^{\text{SoVTP}}_{new}=\left[T_{\text{context}};\,T_{\text{body}};\,T_{\text{AU}};\,E_{\text{others}}\right].$ (16)
The enhanced input $(V^{\text{new}},T^{\text{SoVTP}}_{new})$ is fed into the VLLM $\Theta$. The model analyzes and critiques body language, context, action units, and others’ emotions via chain-of-thought reasoning, then outputs the target emotion $A^{o}$ as:
$A^{o}=\Theta(V^{\text{new}},T^{\text{SoVTP}}_{new}).$ (17)
@@ IV-B Dataset details
DFEW [50] is a large-scale, unconstrained video database specifically designed to advance dynamic facial expression recognition in real-world conditions. It comprises 16,372 clips extracted from over 1,500 high-definition movies, each depicting one of the seven basic emotions—anger, disgust, fear, happiness, neutral, sadness, and surprise—under challenging interferences such as extreme illumination, occlusions, and pose variations. It offer a platform for the development of spatiotemporal deep learning methods for dynamic facial expression recognition in the wild.
IV-C Action units